A motel clerk counts his $1 and $10 bills at the end of the day. He finds that he has a total of 64 bills having a combined monetary value of $199. Find the number of bills of each denomination that he has.
Found 2 solutions by fractalier, addingup:
Answer by fractalier(6550)

Solved this one yesterday...please see my solution...
49 $1 bills and
15 $10 bills

Answer by addingup(3677)

x+y= 64; y= 64-x
1x+10y= 199 Substitute for y:
1x+10(64-x)= 199
1x+640-10x= 199
-9x= -441
x= -441/-9
x= 49
He has 49 $1 bills and 64-49= 15 $10 bills